# Vulnerable Connectivity Caused by Local Communities in Spatial Networks

### v2

@@ Acknowledgements
Supporting Information
S1 Fig. fukuoka_descend_1024_community_2in1.pdf: Visualization of community structures in Fukuoka before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
S2 Fig. fukuoka_ascend_1024_community_2in1.pdf: Visualization of community structures in Fukuoka before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to S1 Fig.
S3 Fig. hiroshima_descend_1024_community_2in1.pdf: Visualization of community structures in Hiroshima before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
S4 Fig. hiroshima_ascend_1024_community_2in1.pdf: Visualization of community structures in Hiroshima before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to S3 Fig.
S5 Fig. keihan_descend_1024_community_2in1.pdf: Visualization of community structures in Keihan before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
S6 Fig. keihan_ascend_1024_community_2in1.pdf: Visualization of community structures in Keihan before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to S5 Fig.
S7 Fig. nagoya_descend_1024_community_2in1.pdf: Visualization of community structures in Nagoya before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
S8 Fig. nagoya_ascend_1024_community_2in1.pdf: Visualization of community structures in Nagoya before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to S7 Fig.
S9 Fig. sendai_descend_1024_community_2in1.pdf: Visualization of community structures in Sendai before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
S10 Fig. sendai_ascend_1024_community_2in1.pdf: Visualization of community structures in Sendai before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to S9 Fig.
S11 Fig. sapporo_descend_1024_community_2in1.pdf: Visualization of community structures in Sapporo before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
S12 Fig. sapporo_ascend_1024_community_2in1.pdf: Visualization of community structures in Sapporo before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to S11 Fig.
S13 Fig. tokyo_rng_qrqc_3in1.pdf: Increasing modularity $Q$ vs. decreasing robustness index $R^{RB}$ or critical fraction ${q_{c}}^{P}{RB}$ for varying the size $N$ in Tokyo RNG networks.
S14 Fig. tokyo_gg_qrqc_3in1.pdf: Increasing modularity $Q$ vs. decreasing robustness index $R^{RB}$ or critical fraction ${q_{c}}^{P}{RB}$ for varying the size $N$ in Tokyo GG networks.
S15 Fig. tokyo_1024_rnqc2grid_2in1.pdf: Scatter plots show relation between network robustness measures (a for $R$ and b for $q_{c}$) and the proportion of grid-like parts against random failures (RF). Networks with N = 1024 nodes are considered, where Pop. networks (green) show notably higher proportions of grid-like parts compared to Inv. (red) and Uni. (blue) networks. See the text at the end of subsection 3.3 for the detail.
S16 Fig. Fukuoka_100_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Fukuoka networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
S17 Fig. Hiroshima_100_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Hiroshima networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
S18 Fig. Keihan_100_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Keihan networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
S19 Fig. Nagoya_100_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Nagoya networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S20 Fig. Tokyo_100_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Tokyo networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
S21 Fig. Sendai_100_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Sendai networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
S22 Fig. Sapporo_100_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Sapporo networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
S23 Fig. Fukuoka_1024_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Fukuoka networks with $N=1024$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
S24 Fig. Hiroshima_1024_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Hiroshima networks with $N=1024$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
S25 Fig. Keihan_1024_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Keihan networks with $N=1024$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
S26 Fig. Nagoya_1024_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Nagoya networks with $N=1024$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
S27 Fig. Sendai_1024_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Sendai networks with $N=1024$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
S28 Fig. Sapporo_1024_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Sapporo networks with $N=1024$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
S29 Fig. Fukuoka_10000_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Fukuoka networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
S30 Fig. Hiroshima_10000_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Hiroshima networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
S31 Fig. Keihan_10000_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Keihan networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
S32 Fig. Nagoya_10000_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Nagoya networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S33 Fig. Tokyo_10000_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Tokyo networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
S34 Fig. Sendai_10000_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Sendai networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
S35 Fig. Sapporo_10000_3combined_4in1.pdf: Robustness against recalculated betweenness (RB) attacks for Sapporo networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
S36 Fig. r2q_100_2in1.pdf Relation between robustness index $R^{RB}$ and modularity $Q$ in networks with $N$ = 100 nodes.
S37 Fig.r2q_10000_2in1.pdf Relation between robustness index $R^{RB}$ and modularity $Q$ in networks with $N$ = 10000 nodes
S6 Table q_values_100: Modularity $Q$ in networks with 100 nodes for seven major Japanese areas. Higher values indicate stronger community structures. Values with upper-triangles ($\triangle$) or lower-triangles ($\triangledown$) indicate where the cases of Pop. and Inv. have higher or lower modularity than the case of Uni. for both RNG and GG. Note the generally higher modularity in Pop. and Inv. compared to Uni. networks for both RNG and GG.
Appendix
.1 Visualization maps. before attacks
@@
S1 Figure: Visualization of community structures in Fukuoka before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
@@
S2 Figure: Visualization of community structures in Fukuoka before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to Figure S1 Figure.
@@
S3 Figure: Visualization of community structures in Hiroshima before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
@@
S4 Figure: Visualization of community structures in Hiroshima before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to Figure S3 Figure.
@@
S5 Figure: Visualization of community structures in Keihan before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
@@
S6 Figure: Visualization of community structures in Keihan before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to Figure S5 Figure.
@@
S7 Figure: Visualization of community structures in Nagoya before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
@@
S8 Figure: Visualization of community structures in Nagoya before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to Figure S7 Figure.
@@
S9 Figure: Visualization of community structures in Sendai before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
@@
S10 Figure: Visualization of community structures in Sendai before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to Figure S9 Figure.
@@
S11 Figure: Visualization of community structures in Sapporo before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
@@
S12 Figure: Visualization of community structures in Sapporo before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to Figure S11 Figure.
3
.2 Visualization maps, after 10% node removals
.3 Q,R and qc relations
@@
S13 Figure: Increasing modularity $Q$ vs. decreasing robustness index $R^{RB}$ or critical fraction $q_{c}^{RB}$ for varying the size $N$ in Tokyo RNG networks.
@@
S14 Figure: Increasing modularity $Q$ vs. decreasing robustness index $R^{RB}$ or critical fraction $q_{c}^{RB}$ for varying the size $N$ in Tokyo GG networks.
.4 scatters of Q vs R or qc
@@
S15 Figure: Scatter plots show relation between network robustness measures (a for $R$ and b for $q_{c}$) and the proportion of grid-like parts against random failures (RF). Networks with $N=1024$ nodes are considered, where Pop. networks (green) show notably higher proportions of grid-like parts compared to Inv. (red) and Uni. (blue) networks. See the text at the end of subsection 3.3 for the detail.
.5 Robustness against RB attacks
@@
S16 Figure: Robustness against recalculated betweenness (RB) attacks for Fukuoka networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S17 Figure: Robustness against recalculated betweenness (RB) attacks for Hiroshima networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S18 Figure: Robustness against recalculated betweenness (RB) attacks for Keihan networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S19 Figure: Robustness against recalculated betweenness (RB) attacks for Nagoya networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S20 Figure: Robustness against recalculated betweenness (RB) attacks for Tokyo networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S21 Figure: Robustness against recalculated betweenness (RB) attacks for Sendai networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S22 Figure: Robustness against recalculated betweenness (RB) attacks for Sapporo networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S23 Figure: Robustness against recalculated betweenness (RB) attacks for Fukuoka networks with $N=1024$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S24 Figure: Robustness against recalculated betweenness (RB) attacks for Hiroshima networks with $N=1024$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S25 Figure: Robustness against recalculated betweenness (RB) attacks for Keihan networks with $N=1024$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.

### v4

@@ Acknowledgements
Fig 1: Illustration of connection constraints for (a) RNG and (b) GG. A link colored by red is established between two nodes colored by blue, when no other node exists within the shaded area.
Fig 2: Coverage diagram of radio waves between two base stations in wireless communication. The ranges of strong beams are shown by blue and orange shades. The center circle represents the signal interference area. If other base stations exist within it, the two stations cannot be connected.
Fig 3: Visualization of community structure in (left) RNG and (right) GG (a)(b) as the original and (c)(d) after 10% nodes removal for Tokyo with $N=1024$ nodes located decreasing order of population (Pop.). Note that RNG is a subgraph of GG. Different communities detected by the Louvain method are colored by red, yellow, green, and blue. Each figure depicts a spatial network on a two-dimensional Euclidean plane, where node’s locations correspond to the centroid locations of 500m × 500m population mesh blocks. The horizontal and vertical axes represent planar distances derived from these centroid positions.
Fig 4: Visualization of community structure in (left) RNG and (right) GG (a)(b)as the original and (c)(d) after 10% nodes removal for Tokyo with $N=1024$ nodes located inverse decreasing order of population (Inv.). Note that RNG is a subgraph of GG. Different communities detected by the Louvain method are colored by red, yellow, green, and blue. Each figure depicts a spatial network on a two-dimensional Euclidean plane, where node’s locations correspond to the centroid locations of 500m × 500m population mesh blocks. The horizontal and vertical axes represent planar distances derived from these centroid positions.
Fig 5: Relation between modularity $Q$ and robustness index $R^{RB}$ in networks with $N=1024$ nodes of 7 Japanese areas against RB attacks. Each point represents the values of $R^{RB}$ and $Q$ for the original (triangles) and 2D lattice (crosses) networks with node’s locations based on Pop. (green/yellow), Inv. (red/magenta), and Uni. (blue/cyan). Links are created by RNG (a) or GG (b). Inset in (a) enlarges the lower-right area with densely points. A monotone decreasing is observed: networks with stronger communities (higher $Q$) tend to have lower robustness index $R^{RB}$. Similar relations are observed in RNG, while the ranges of $Q$ and $R^{RB}$ differ slightly.
Fig 6: Degree distributions $P(k)$ in the original Tokyo networks with $N=1024$ and with node’s locations based on Pop., Inv., and Uni.. All cases have bell-shaped forms with peaks around $k=2$ to $4$.
Fig 7: Relation between robustness index $R^{RB}$ and sparsity index $SI(G_{w})$ for networks with $N=1024$ nodes in seven Japanese areas. Each point represents the results for (a) RNG and (b) GG with node’s locations based on Pop. (green), Inv. (red), and Uni. (blue). Colored points show clear monotone decreasing that networks with higher sparsity index $SI(G_{w})$ tend to have lower robustness index $R^{RB}$. We note that the robustness index $R^{RB}$ on the vertical axis varies within a narrow range, while the sparsity index $SI(G_{w})$ on the horizontal axis spans a much wider range.
Fig 8: Relation between sparsity index $SI(G_{w})$ and modularity $Q$ for the networks with $N=1024$ nodes in seven Japanese areas. Each point represents the result for (a) RNG and (b) GG with node’s locations based on Pop. (green triangles), Inv. (red squares), or Uni. (blue circle). Colored points show clear monotone increasing: networks with higher spatial sparsity index $SI(G_{w})$ tend to have higher $Q$. Pop.-based networks (green triangles) have the highest $SI$ values because of long-range link between spatially concentrated urban regions and peripheral regions, while Uni.-based (blue circle) has the lowest $SI$ because of uniformly positioned node’s locations. We note that the scales of vertical and horizontal axes are quite different. In other words, the vertical differences are ignorable.
@@
Fig 9: Robustness against recalculated betweenness (RB) attacks in Tokyo networks with $N=1024$ nodes. For rewired (randomized networks) lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) the relative size $S^{1st}(q)/N$ of largest connected component, and (c) (d) the critical fraction $q_{c}$ at the peak of the relative size $S^{2nd}(q)/N$ of second largest component.
@@
Fig 10: Robustness against recalculated betweenness (RB) attacks in Tokyo networks with $N=1024$ nodes. For 2DL (relocated networks) lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) the relative size $S^{1st}(q)/N$ of largest connected component, and (c) (d) the critical fraction $q_{c}$ at the peak of the relative size $S^{2nd}(q)/N$ of second largest component.
@@
S1 Fig: Visualization of community structures in Fukuoka before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
@@
S2 Fig: Visualization of community structures in Fukuoka before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to Figure S1 Fig.
@@
S3 Fig: Visualization of community structures in Hiroshima before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
@@
S4 Fig: Visualization of community structures in Hiroshima before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to Figure S3 Fig.
@@
S5 Fig: Visualization of community structures in Keihan before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
@@
S6 Fig: Visualization of community structures in Keihan before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to Figure S5 Fig.
@@
S7 Fig: Visualization of community structures in Nagoya before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
@@
S8 Fig: Visualization of community structures in Nagoya before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to Figure S7 Fig.
@@
S9 Fig: Visualization of community structures in Sendai before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
@@
S10 Fig: Visualization of community structures in Sendai before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to Figure S9 Fig.
@@
S11 Fig: Visualization of community structures in Sapporo before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
@@
S12 Fig: Visualization of community structures in Sapporo before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to Figure S11 Fig.
@@
S13 Fig: Increasing modularity $Q$ vs. decreasing robustness index $R^{RB}$ or critical fraction $q_{c}^{RB}$ for varying the size $N$ in Tokyo RNG networks.
@@
S14 Fig: Increasing modularity $Q$ vs. decreasing robustness index $R^{RB}$ or critical fraction $q_{c}^{RB}$ for varying the size $N$ in Tokyo GG networks.
@@
S15 Fig: Scatter plots show relation between network robustness measures (a for $R$ and b for $q_{c}$) and the proportion of grid-like parts against random failures (RF). Networks with $N=1024$ nodes are considered, where Pop. networks (green) show notably higher proportions of grid-like parts compared to Inv. (red) and Uni. (blue) networks. See the text at the end of subsection 3.3 for the detail.
@@
S16 Fig: Robustness against recalculated betweenness (RB) attacks for Fukuoka networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S17 Fig: Robustness against recalculated betweenness (RB) attacks for Hiroshima networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S18 Fig: Robustness against recalculated betweenness (RB) attacks for Keihan networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S19 Fig: Robustness against recalculated betweenness (RB) attacks for Nagoya networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S20 Fig: Robustness against recalculated betweenness (RB) attacks for Tokyo networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S21 Fig: Robustness against recalculated betweenness (RB) attacks for Sendai networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S22 Fig: Robustness against recalculated betweenness (RB) attacks for Sapporo networks with $N=100$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S23 Fig: Robustness against recalculated betweenness (RB) attacks for Fukuoka networks with $N=1024$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S24 Fig: Robustness against recalculated betweenness (RB) attacks for Hiroshima networks with $N=1024$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S25 Fig: Robustness against recalculated betweenness (RB) attacks for Keihan networks with $N=1024$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.